OpenStudy (anonymous):

Evaluate: 7/8 - 3/4 times 1/4

OpenStudy (abmon98):

Use PEDMAS P Parentheses first E Exponents (ie Powers and Square Roots, etc.) MD Multiplication and Division (left-to-right) AS Addition and Subtraction (left-to-right)

OpenStudy (anonymous):

This is what I have so far \[\frac{ 7 }{ 8 } - \frac{ 3 }{ 4 } \times \frac{ 1 }{ 4 } = \frac{ 7 }{ 8} - \frac{ 3 }{ 16 }\] I don't understand how to do the LCD of 16, that's where I keep getting stumped. @Abmon98

OpenStudy (anonymous):

I also just noticed I posted this in the wrong section fml. I'm sorry!